OpenStudy (anonymous):

A speed (L/T) divided by a force (ML/T2) yields?

OpenStudy (anonymous):

is it T/M^-1?

OpenStudy (mathteacher1729):

Just to be clear: L = Length T = Time M = Mass So a speed divided by a force is: \[\frac{\frac{L}{T}}{\frac{ML}{T^2}}=\frac{L}{T}\frac{T^2}{ML}\] I obtained the right hand side by the rule of dividing fractions ("invert the denominator and multiply by the numerator"). Now we cancel factors...